The data below show real wage levels at various percentiles of the wage distribution for material-handling labourers in Fiji

Percentile 1975 1995
10th $4.26 $2.97
50th 6.02 3.89
90th 10.44 9.07

A. Calculate w90/w50, w50/w10, and w90/w10 percentile wage differentials for both years.
B. Describe the changes in wage inequality in this local market
C. Use the relative demand and supply model to explain the effect of increased immigrants into this labour market on relative wage and employment outcomes
1
Expert's answer
2019-04-18T09:35:47-0400

A. In 1975: w90/w50 = 10.44/6.02 = 1.73,

w50/w10 = 1.41,

w90/w10 = 2.45.

In 1995: w90/w50 = 9.07/3.89 = 2.33,

w50/w10 = 1.31,

w90/w10 = 3.05.

B. The wage inequality increased in this local market.

C. Increased immigrants into this labour market will increase supply of labour, so relative wage will decrease and employment will increase.